the perimeter of a basketball court is 228 ft. If the length of the court is 6 ft longer than twice the the width. find the dimensions of the basketball court. Please include variables and representation.
Answer by oberobic(2304)

P = perimeter = 2(Length + Width)
.
L = 2W +6
.
P = 228 ft
.
substitute for L
.
2( 2W+6 +W ) = 228
.
4W +12 +2W = 228
.
6W = 216
.
W = 36
.
L = 2W +6= 72+6 = 78
.
Check perimeter
.
2(36) + 2(78) = 72 + 156 = 228
Correct
.
Answer: Length = 78; Width = 36
.
Done.